### v1

# Current affiliation: ]Fakultät für Physik, Technische Universität Dortmund, 44227 Dortmund, Germany Spin Liquid Mimicry in the Hydroxide Double Perovskite CuSn(OD)6 Induced by Correlated Proton Disorder

Abstract
The face-centered-cubic lattice is composed of edge-sharing tetrahedra, making it a leading candidate host for strongly frustrated magnetism, but relatively few face-centered frustrated materials have been investigated. In the hydroxide double perovskite CuSn(OH)6, magnetic frustration of the Cu2+ quantum spins is partially relieved by strong Jahn-Teller distortions. Nevertheless, the system shows no signs of long-range magnetic order down to 45 mK and instead exhibits broad thermodynamic anomalies in specific heat and magnetization, indicating short-range dynamical spin correlations — a behavior typical of quantum spin liquids. We propose that such an unusual robustness of the spin-liquid-like state is a combined effect of quantum fluctuations of the quantum spins , residual frustration on the highly distorted face-centered Cu2+ sublattice, and correlated proton disorder. Similar to the disorder-induced spin-liquid mimicry in YbMgGaO4 and herbertsmithite, proton disorder destabilizes the long-range magnetic order by introducing randomness into the magnetic exchange interaction network. However, unlike the quenched substitutional disorder on the magnetic sublattice, which is difficult to control, proton disorder can in principle be tuned through pressure-driven proton ordering transitions. This opens up the prospect of tuning the degree of disorder in a magnetic system to better understand its influence on the magnetic ground state.
I Introduction
A quantum spin liquid (QSL), proposed by Anderson in 1973 [1], is an exotic state of magnetic matter that exhibits liquid-like behavior with strong quantum fluctuations and avoids long-range magnetic order even at absolute zero temperature [2, 3, 4]. QSLs are both fundamentally intriguing and a promising platform for quantum computing, and extensive theoretical and experimental efforts have been made to classify different types of QSLs and to find their material realizations [5, 6, 7, 8, 9, 10, 11, 12, 13]. Avoiding long-range order despite strong interactions among the magnetic moments is generally accomplished through frustration, whereby exchange interactions compete with each other, preventing the system from readily selecting a unique ground state. Most spin liquids are highly fragile, as their stability is confined to a tiny region of parameter space. In the best-known candidate spin-liquid systems, minute lattice distortions or other deviations from the idealized spin model typically stabilize spin order at sufficiently low temperatures [14, 15, 16].
The identification of spin-liquid states is challenging due to the lack of a discernible order parameter. An absence of magnetic order can be attributed to various factors, including quenched substitutional disorder or site intermixing. These phenomena show comparable signatures to spin liquids in thermodynamic measurements. This particular form of behavior, in which a QSL-like state is induced by quenched atomic disorder, has been termed “spin-liquid mimicry” to distinguish it from the true spin-liquid ground states, which persist even within an idealized, perfectly spatially periodic system [17]. Notable examples of such behavior are found in the mineral herbertsmithite ZnCu3(OH)6Cl2 [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] and the triangular-lattice compound YbMgGaO4 [32, 33, 17].
Hydroxide perovskites with correlated proton disorder [34, 35, 36, 37] appear promising in this respect, because their hydrogen sites obey analogies of the ice rules known from water ice and spin ices, and it is expected that this correlated proton disorder can be tuned by hydrostatic pressure [38] via a sequence of proton ordering transitions, similar to those known in conventional water ice [39, 40, 41]. In this paper we focus on CuSn(OH)6, a hydroxide double perovskite crystallizing in the orthorhombic space group [38]. The crystal structure of synthetic deuterated mushistonite, which we refined previously [38] from a combination of neutron and x-ray diffraction, is shown in Fig. 1. It consists of alternating [Cu2+(OD)6] and [Sn4+(OD)6] octahedra, with significant Jahn-Teller distortion of the Cu environment.
In this work, we investigate the low-temperature magnetic properties of mushistonite as the first example of a material exhibiting spin-liquid mimicry induced by proton disorder. We investigate its magnetic behavior through heat capacity, magnetic susceptibility, and powder neutron diffraction measurements, and find that it is likely a uniquely suitable material for continuously tuning spin-liquid mimicry.
II Experimental details
II.1 Sample preparation
Phase-pure powder samples of CuSn(OH)6 were synthesized by co-precipitation from aqueous solutions of CuClH2O and Na2Sn(OH)6 at room temperature. Equimolar amounts of the starting materials were ground together before dissolving them in water with concentration 0.16 mol/L. After six days of vigorous stirring the precipitate was filtered, washed with ethanol, and dried in a vacuum drying chamber. The deuterated version, CuSn(OD)6, was synthesized by the same procedure from a solution in D2O. The resulting pale blue powder contains intergrown cubic crystallites less than 1 m in size.
II.2 Specific heat and magnetization
Low-temperature specific-heat measurements were performed by the relaxation time method on a thin cold-pressed pellet of CuSn(OH)6 using a Quantum Design Physical Property Measurement System (PPMS) both using a standard 4He specific heat puck and a respective puck in a 3He refrigerator. The sample was mounted to the measurement platform using Apiezon N grease; contributions from the sample holder and grease were subtracted. Multiple data points were collected at each temperature and averaged.
Magnetization measurements down to 2 K were performed by vibrating sample magnetometry (VSM) in a Cryogenic Ltd. Cryogen-Free Measurement System (CFMS), under zero-field-cooled-warming and field-cooled conditions. Four-quadrant – loops were measured to 14 T at several temperatures. Samples were loaded inside gel capsules which were inserted into plastic straws. Additional magnetization measurements down to 0.45 K were performed in a Quantum Design Magnetic Property Measurement System (MPMS-XL) equipped with an iHelium3 3He refrigerator insert.
High-field magnetization measurements up to 35 T at 1.4 K were conducted at the Hochfeld-Magnetlabor Dresden (HLD), Helmholtz-Zentrum Dresden-Rossendorf (HZDR), in Dresden, Germany, using a pulsed magnet with a rise time of 7 ms and a total pulse duration of 25 ms. The magnetization was obtained by integrating the voltage induced in a compensated coil system surrounding the sample [43].
II.3 Neutron powder diffraction
Neutron powder diffraction (NPD) measurements were conducted on CuSn(OD)6 using the HRPT diffractometer [44] at SINQ, PSI, in Villigen, Switzerland, using neutrons of wavelength 2.450 Å. The counting times were approximately 26 hours each at 1.6 and 6 K.
Additional NPD measurements were performed with 2.526-Å neutrons on the D1B diffractometer [45] at the Institut Laue-Langevin (ILL) in Grenoble, France, using a dilution refrigerator. The counting times were approximately 8 hours at 0.05 K and 2 hours at 10 K. The wavelength was selected using the (002) reflection from pyrolytic graphite.
III Results and Discussion
The specific heat of CuSn(OH)6 is shown in Fig. 2 (a). Based on its color the material is assumed to be electrically insulating, so this can be described as . The lattice contribution is typically modeled using Debye, Einstein, or a combination of both terms. However, in this case, a phenomenological function including two Debye and two Einstein terms is used:
The contribution of the Debye heat capacity terms is
where is the number of participating atoms in the solid, is the Boltzmann constant and is the Debye temperature corresponding to the acoustic phonon vibrations in the sample. The contribution of the Einstein heat capacity terms at a constant volume is
where is the Einstein temperature corresponding to the optical phonons.
Figure 2 (a) shows results of the fit in the temperature range of 20–210 K using the single and double Debye-Einstein models, i.e., with two and four contributions, respectively. The double Debye-Einstein model shows significantly better agreement with the experimental data.
The double Debye-Einstein model fit was subtracted from the specific-heat data over the entire temperature range to obtain the magnetic contribution to the specific heat , shown in Fig. 2 (b). The broad magnetic hump at 4.1 K suggests the presence of short-range correlations. The low-temperature behavior follows neither a power-law dependence as with or 2, characteristic of gapless QSLs, nor an exponential as ), characteristic of gapped QSLs. The hump also cannot be described as a Schottky-like anomaly. This feature broadens and shifts to lower temperature when a magnetic field is applied. The plot of in the inset of Fig. 2 (b) shows an additional magnetic-field-sensitive anomaly at about 0.8 K, which is barely visible in . The entropy in this peak is extremely small. As shown below, this transition does not seem to be associated with long-range order.
The magnetic entropy obtained by integrating is shown in Fig. 2 (c). A point at (0,0) was added to all curves for an approximate extrapolation to ; this extrapolated region accounts for 0.13 J mol-1 K-1. The magnetic entropy saturates around 30 K, far above the specific-heat anomalies as expected for frustrated spin systems, but it asymptotes only 85% of . This suggests that not all spin degrees of freedom are available at low temperatures. This may indicate a magnetic transition at temperatures lower than our minimum of 0.45 K, a model subtraction issue, or quantum entanglement suppressing classical thermal excitations.
As shown in Fig. 2 (d), the magnetic susceptibility data also show a broad anomaly, which appears as a broad peak in the first derivative with a maximum at 4.1 K. No further significant features are observed down to 0.45 K in low field, as shown in Fig. 2 (e). A subtle crossover appears in our magnetization data collected at 1 T, where a slight kink appears at approximately 1.3 K. However, due to the small magnitude of these changes, it is challenging to determine the nature of this feature. Fig. 2(e) shows both field-cooled and zero-field-cooled magnetization data, and no significant differences are found between these, indicating the absence of spin-glass behavior or any freezing of the spins. The Curie-Weiss temperature was obtained from a fit of as shown in the inset of Fig. 2 (e). The estimated Curie-Weiss temperature K, indicating dominant antiferromagnetic interactions, while the paramagnetic moment is , slightly higher than the expected value for . The apparent lack of long-range magnetic order down to temperatures more than an order of magnitude below the Curie-Weiss temperature suggests strong frustration.
Fig. 2 (f) shows the magnetic field dependence of the magnetization at several temperatures, and up to 35 T at 1.4 K. An apparent saturation near 1 per Cu is reached at roughly 18 T. The 18-T energy scale exceeds that of the Curie-Weiss temperature by nearly a factor of two, suggesting a complex frustrated network of significant ferromagnetic and antiferromagnetic interactions which conspire to reduce the Curie-Weiss temperature.
To check for signs of magnetic order across the potential transitions at 0.8 and 4.1 K, we performed NPD measurements at 0.05 and 10 K at D1B, ILL, and at 1.6 and 6.0 K on HRPT, at PSI (see Fig. 3). We have already reported the crystal structure in detail elsewhere [38]. No features associated with magnetic long-range — or even short-range — order are seen down to 50 mK, and no significant change in the background was observed at any angle. This is in stark contrast to the sister compound MnSn(OD), in which broad magnetic peaks were observed at low temperatures, associated with propagation vectors of and/or but having a correlation length of only a few unit cells [46]. We note that the emergence of order on the hydrogen site would be expected to leave strong signatures in the neutron diffraction pattern, so this can also be excluded.
Taking all our data together, we see no evidence of even short-range magnetic order down to 50 mK, despite a Curie-Weiss temperature of 7.1 K and an apparent interaction energy scale of 18 T. The nearly 2.5 orders of magnitude separating the temperature of our neutron measurement from the interaction energy scale suggests very strong frustration in CuSn(OD)6. However, the suppression of magnetic order may also arise from quantum fluctuations. The Mn compound has a similar but higher-symmetric crystal structure with analogous proton disorder, but its magnetic moments are classical spins which are not expected to experience significant quantum fluctuations. The lower symmetry in the Cu analog would ordinarily be expected to partially relieve frustration, making long-range order more stable, so the lack of order suggests that quantum fluctuations play a key role in CuSn(OD)6.
Having established that CuSn(OD)6 does not order down to 50 mK, the question is what it does instead. One possibility is that it is a quantum spin liquid. We see evidence that quantum fluctuations may play an important role in the material, which would argue in favor of this scenario, but the strong Jahn-Teller distortions of the Cu coordination sphere lead to a rather low-symmetry crystal structure with a large number of inequivalent interactions, which would be expected to break the degeneracy required for a QSL ground state. It is conceivable, although extremely unlikely, that these interactions are tuned to hit a pocket of QSL by sheer coincidence. However, the correlated disorder on the hydrogen site adds another random fluctuation to the exchanges. This makes a QSL extremely unlikely, in our opinion. Instead, we propose that CuSn(OD)6 exhibits spin liquid mimicry.
The significant Zn/Cu cross-substitution in herbertsmithite [18, 19, 20, 21, 22] leads to similar physics. No long-range magnetic order is observed in herbertsmithite down to 50 mK despite strong antiferromagnetic interactions [23, 24, 25, 26]; however, the nature of the spin-liquid ground state remains under debate, and neutron and NMR data provide conflicting evidence. Some studies report a gapped [22, 27, 28] ground state while others suggest a gapless [24, 29, 30] one. Similarly, triangular-lattice YbMgGaO4 with [32] shows no magnetic order down to 50 mK, SR measurements indicate persistent spin dynamics compatible with a U(1)-type QSL state, and inelastic neutron scattering further revealed a gapless continuum of magnetic excitations at 60 mK [33]. However, random site mixing between Mg2+ and Ga3+ leads to a locally disordered crystalline environment around Yb3+ ions, resulting in spatially varying exchange interactions [17]. This structural randomness mimics spin-liquid behavior in YbMgGaO4.
In both herbertsmithite and YbMgGaO4, spin-liquid mimicry arises from quenched substitutional disorder that is difficult to control. In CuSn(OD)6, however, the disorder arises from ice-rule physics among the hydrogen atoms which fill the perovskite A site. In water ice this can be tuned with pressure, leading to a rich phase diagram, and our previous neutron diffraction pressure study indicated that the hydrogen sites in CuSn(OD)6 are likely to merge at pressures on the order of 20 GPa [38]. This is an experimentally-accessible pressure. It should be possible to continuously tune the correlated hydrogen disorder out of existence while performing a variety of measurements, replacing the spin-liquid-like ground state we observe with either magnetic order or a quantum spin liquid.
With magnetic order excluded as an explanation, and the ground state likely a spin liquid mimic, the hump in the specific heat at 4.1 K may indicate gapless spin-liquid mimicry behavior, while the low-temperature anomaly at 0.8 K may indicate the emergence of a spin gap. This would suggest that temperature-dependent magnetic exchange interactions could modify the excitation spectrum, leading to a crossover from gapless to possibly gapped spin-liquid mimicry.
IV Conclusions
CuSn(OH)6 does not order magnetically down to 50 mK, and we find no evidence of spin freezing down to at least 0.45 K. Despite significant antiferromagnetic exchange interactions, a combination of quantum fluctuations, frustration, and hydrogen disorder prevents the formation of long-range magnetic order or any other static spin arrangement. Given the strong hydrogen disorder, we propose that CuSn(OH)6 is a rare example of spin-liquid mimicry, driven in this case by tunable structural disorder in the proton sublattice. In a unique difference from other such systems, however, in CuSn(OH)6 this disorder can be tuned. The ability to control hydrogen disorder via external pressure [38] makes CuSn(OH)6 (and possibly some other hydroxide perovskites) a particularly promising platform for probing the interplay between quantum magnetism and structural disorder and exploring spin-liquid mimicry as a function of disorder strength.
Data Availability
Samples and data are available upon reasonable request from D. C. Peets or D. S. Inosov; data underpinning this work is available from Refs. 47, 48.
Acknowledgements.
We gratefully acknowledge L. Zviagina for help with the pulsed-field magnetization experiments. This project was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through: individual grants IN 209/12-1, DO 590/11-1 (Project No. 536621965), and PE 3318/2-1 (Project No. 452541981); through projects B01, B03, C01, and C03 of the Collaborative Research Center SFB 1143 (Project No. 247310070); and through the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Materials — ct.qmat (EXC 2147, Project No. 390858490). The authors acknowledge the support of the Institut Laue-Langevin, Grenoble, France, as well as the HLD at HZDR, member of the European Magnetic Field Laboratory (EMFL). This work is based in part on experiments performed at the Swiss spallation neutron source SINQ, Paul Scherrer Institute, Villigen, Switzerland.References
- Anderson [1973] P. Anderson, Resonating valence bonds: A new kind of insulator?, Mater. Res. Bull. 8, 153 (1973).
- Savary and Balents [2016] L. Savary and L. Balents, Quantum spin liquids: a review, Rep. Prog. Phys. 80, 016502 (2016).
- Balents [2010] L. Balents, Spin liquids in frustrated magnets, Nat. 464, 199 (2010).
- Ramirez and Syzranov [2025] A. P. Ramirez and S. V. Syzranov, Short-range order and hidden energy scale in geometrically frustrated magnets, Mater. Adv. 6, 1213 (2025).
- Mei et al. [2017] J.-W. Mei, J.-Y. Chen, H. He, and X.-G. Wen, Gapped spin liquid with topological order for the kagome Heisenberg model, Phys. Rev. B 95, 235107 (2017).
- Lu and Lu [2024] S. Lu and Y.-M. Lu, Detecting symmetry fractionalization in gapped quantum spin liquids by magnetic impurities, Phys. Rev. B 110, L100401 (2024).
- Miksch et al. [2021] B. Miksch, A. Pustogow, M. J. Rahim, A. A. Bardin, K. Kanoda, J. A. Schlueter, R. Hübner, M. Scheffler, and M. Dressel, Gapped magnetic ground state in quantum spin liquid candidate -(BEDT-TTF)2Cu2(CN)3, Sci. 372, 276 (2021).
- Feng et al. [2017] Z. Feng, Z. Li, X. Meng, W. Yi, Y. Wei, J. Zhang, Y.-C. Wang, W. Jiang, Z. Liu, S. Li, F. Liu, J. Luo, S. Li, G.-q. Zheng, Z. Y. Meng, J.-W. Mei, and Y. Shi, Gapped spin-1/2 spinon excitations in a new kagome quantum spin liquid compound CuZn(OH)FBr, Chin. Phys. Lett. 34, 077502 (2017).
- Yamashita et al. [2011] S. Yamashita, T. Yamamoto, Y. Nakazawa, M. Tamura, and R. Kato, Gapless spin liquid of an organic triangular compound evidenced by thermodynamic measurements, Nat. Commun. 2, 275 (2011).
- Liu et al. [2022a] J. Liu, L. Yuan, X. Li, B. Li, K. Zhao, H. Liao, and Y. Li, Gapless spin liquid behavior in a kagome Heisenberg antiferromagnet with randomly distributed hexagons of alternate bonds, Phys. Rev. B 105, 024418 (2022a).
- Fujihala et al. [2020] M. Fujihala, K. Morita, R. Mole, S. Mitsuda, T. Tohyama, S.-I. Yano, D. Yu, S. Sota, T. Kuwai, A. Koda, H. Okabe, H. Lee, S. Itoh, T. Hawai, T. Masuda, H. Sagayama, A. Matsuo, K. Kindo, S. Ohira-Kawamura, and K. Nakajima, Gapless spin liquid in a square-kagome lattice antiferromagnet, Nat. Commun. 11, 3429 (2020).
- Barkeshli et al. [2013] M. Barkeshli, H. Yao, and S. A. Kivelson, Gapless spin liquids: Stability and possible experimental relevance, Phys. Rev. B 87, 140402 (2013).
- Liu et al. [2022b] W.-Y. Liu, J. Hasik, S.-S. Gong, D. Poilblanc, W.-Q. Chen, and Z.-C. Gu, Emergence of gapless quantum spin liquid from deconfined quantum critical point, Phys. Rev. X 12, 031039 (2022b).
- Kulbakov et al. [2021] A. A. Kulbakov, S. M. Avdoshenko, I. Puente-Orench, M. Deeb, M. Doerr, P. Schlender, T. Doert, and D. S. Inosov, Stripe- magnetic order in the triangular-lattice antiferromagnet KCeS2, J. Condens. Matter Phys. 33, 425802 (2021).
- Avdoshenko et al. [2022] S. M. Avdoshenko, A. A. Kulbakov, E. Häußler, P. Schlender, T. Doert, J. Ollivier, and D. S. Inosov, Spin-wave dynamics in the KCeS2 delafossite: A theoretical description of powder inelastic neutron-scattering data, Phys. Rev. B 106, 214431 (2022).
- Xie et al. [2023] T. Xie, A. A. Eberharter, J. Xing, S. Nishimoto, M. Brando, P. Khanenko, J. Sichelschmidt, A. A. Turrini, D. G. Mazzone, P. G. Naumov, L. D. Sanjeewa, N. Harrison, A. S. Sefat, B. Normand, A. M. Läuchli, A. Podlesnyak, and S. E. Nikitin, Complete field-induced spectral response of the spin-1/2 triangular-lattice antiferromagnet CsYbSe2, npj Quantum Mater. 8, 48 (2023).
- Zhu et al. [2017] Z. Zhu, P. A. Maksimov, S. R. White, and A. L. Chernyshev, Disorder-induced mimicry of a spin liquid in YbMgGaO4, Phys. Rev. Lett. 119, 157201 (2017).
- Freedman et al. [2010] D. E. Freedman, T. H. Han, A. Prodi, P. Müller, Q.-Z. Huang, Y.-S. Chen, S. M. Webb, Y. S. Lee, T. M. McQueen, and D. G. Nocera, Site specific x-ray anomalous dispersion of the geometrically frustrated kagomé magnet, herbertsmithite, ZnCu3(OH)6Cl2, J. Am. Chem. Soc. 132, 16185 (2010).
- Zorko et al. [2017] A. Zorko, M. Herak, M. Gomilšek, J. van Tol, M. Velázquez, P. Khuntia, F. Bert, and P. Mendels, Symmetry reduction in the quantum kagome antiferromagnet herbertsmithite, Phys. Rev. Lett. 118, 017202 (2017).
- Smaha et al. [2020] R. W. Smaha, I. Boukahil, C. J. Titus, J. M. Jiang, J. P. Sheckelton, W. He, J. Wen, J. Vinson, S. G. Wang, Y.-S. Chen, S. J. Teat, T. P. Devereaux, C. Das Pemmaraju, and Y. S. Lee, Site-specific structure at multiple length scales in kagome quantum spin liquid candidates, Phys. Rev. Mater. 4, 124406 (2020).
- Huang et al. [2021] Y. Y. Huang, Y. Xu, L. Wang, C. C. Zhao, C. P. Tu, J. M. Ni, L. S. Wang, B. L. Pan, Y. Fu, Z. Hao, C. Liu, J.-W. Mei, and S. Y. Li, Heat transport in herbertsmithite: Can a quantum spin liquid survive disorder?, Phys. Rev. Lett. 127, 267202 (2021).
- Imai et al. [2011] T. Imai, M. Fu, T. H. Han, and Y. S. Lee, Local spin susceptibility of the kagome lattice in ZnCu3(OD)6Cl2, Phys. Rev. B 84, 020411 (2011).
- Mendels et al. [2007] P. Mendels, F. Bert, M. A. de Vries, A. Olariu, A. Harrison, F. Duc, J. C. Trombe, J. S. Lord, A. Amato, and C. Baines, Quantum magnetism in the paratacamite family: Towards an ideal kagomé lattice, Phys. Rev. Lett. 98, 077204 (2007).
- Helton et al. [2007] J. S. Helton, K. Matan, M. P. Shores, E. A. Nytko, B. M. Bartlett, Y. Yoshida, Y. Takano, A. Suslov, Y. Qiu, J.-H. Chung, D. G. Nocera, and Y. S. Lee, Spin dynamics of the spin- kagome lattice antiferromagnet ZnCu3(OH)6Cl2, Phys. Rev. Lett. 98, 107204 (2007).
- Misguich and Sindzingre [2007] G. Misguich and P. Sindzingre, Magnetic susceptibility and specific heat of the spin- Heisenberg model on the kagome lattice and experimental data on ZnCu3(OH)6Cl2, Eur. Phys. J. B 59, 305 (2007).
- Chitra and Rozenberg [2008] R. Chitra and M. J. Rozenberg, Impurity effects in the quantum kagome system ZnCu3(OH)6Cl2, Phys. Rev. B 77, 052407 (2008).
- Fu et al. [2015] M. Fu, T. Imai, T.-H. Han, and Y. S. Lee, Evidence for a gapped spin-liquid ground state in a kagome Heisenberg antiferromagnet, Science 350, 655 (2015).
- Han et al. [2016] T.-H. Han, M. R. Norman, J.-J. Wen, J. A. Rodriguez-Rivera, J. S. Helton, C. Broholm, and Y. S. Lee, Correlated impurities and intrinsic spin-liquid physics in the kagome material herbertsmithite, Phys. Rev. B 94, 060409 (2016).
- Khuntia et al. [2020] P. Khuntia, M. Velazquez, Q. Barthélemy, F. Bert, E. Kermarrec, A. Legros, B. Bernu, L. Messio, A. Zorko, and P. Mendels, Gapless ground state in the archetypal quantum kagome antiferromagnet ZnCu3(OH)6Cl2, Nat. Phys. 16, 469 (2020).
- de Vries et al. [2008] M. A. de Vries, K. V. Kamenev, W. A. Kockelmann, J. Sanchez-Benitez, and A. Harrison, Magnetic ground state of an experimental kagome antiferromagnet, Phys. Rev. Lett. 100, 157205 (2008).
- Norman [2016] M. R. Norman, : Herbertsmithite and the search for the quantum spin liquid, Rev. Mod. Phys. 88, 041002 (2016).
- Li et al. [2016] Y. Li, D. Adroja, P. K. Biswas, P. J. Baker, Q. Zhang, J. Liu, A. A. Tsirlin, P. Gegenwart, and Q. Zhang, Muon spin relaxation evidence for the U(1) quantum spin-liquid ground state in the triangular antiferromagnet YbMgGaO4, Phys. Rev. Lett. 117, 097201 (2016).
- Li [2019] Y. Li, YbMgGaO4: A triangular-lattice quantum spin liquid candidate, Adv. Quantum Technol. 2, 1900089 (2019).
- Welch and Najorka [2025] M. D. Welch and J. Najorka, Hydroxyperovskites: An overlooked class of potential functional materials, Crystals 15, 251 (2025).
- Lafuente et al. [2015] B. Lafuente, H. Yang, and R. T. Downs, Crystal structure of tetrawickmanite, Mn2+Sn4+(OH)6, Acta Crystallogr. E 71, 234 (2015).
- Kampf et al. [2024] A. R. Kampf, J. Désor, M. D. Welch, C. Ma, and G. Möhn, Zincostottite, ZnGe(OH)6, the zinc analogue of stottite from Tsumeb, Namibia, Mineral. Mag. 88, 1 (2024).
- [37] L. C. Basciano, R. C. Peterson, P. L. Roeder, and I. Swainson, Description of schoenfliesite, MgSn(OH)6, and roxbyite, Cu1.72S, from a 1375 BC shipwreck, Rietveld neutron-diffraction refinement of synthetic schoenfliesite, wickmanite, MnSn(OH)6, and burtite, CaSn(OH)6, Canad. Mineral. , 1203 (1998).
- Kulbakov et al. [2025a] A. A. Kulbakov, E. Häußler, K. K. Parui, A. Mannathanath Chakkingal, N. S. Pavlovskii, V. Y. Pomjakushin, L. Cañadillas Delgado, T. Hansen, D. C. Peets, T. Doert, and D. S. Inosov, Correlated proton disorder in the crystal structure of the double hydroxide perovskite CuSn(OH)6, Phys. Rev. Mater. 9, 033603 (2025a).
- Bernal and Fowler [1933] J. D. Bernal and R. H. Fowler, A theory of water and ionic solution, with particular reference to hydrogen and hydroxyl ions, J. Chem. Phys. 1, 515 (1933).
- Pauling [1935] L. Pauling, The structure and entropy of ice and of other crystals with some randomness of atomic arrangement, J. Am. Chem. Soc. 57, 2680 (1935).
- Malenkov [2009] G. Malenkov, Liquid water and ices: understanding the structure and physical properties, J. Phys.: Condens. Matter 21, 283101 (2009).
- Momma and Izumi [2011] K. Momma and F. Izumi, Vesta 3 for three-dimensional visualization of crystal, volumetric and morphology data, J. Appl. Crystallogr. 44, 1272 (2011).
- Skourski et al. [2011] Y. Skourski, M. D. Kuz’min, K. P. Skokov, A. V. Andreev, and J. Wosnitza, High-field magnetization of Ho2Fe17, Phys. Rev. B 83, 214420 (2011).
- Fischer et al. [2000] P. Fischer, G. Frey, M. Koch, M. Könnecke, V. Pomjakushin, J. Schefer, R. Thut, N. Schlumpf, R. Bürge, U. Greuter, S. Bondt, and E. Berruyer, High-resolution powder diffractometer HRPT for thermal neutrons at SINQ, Physica B Condens. Matter 276–278, 146 (2000).
- Puente Orench et al. [2014] I. Puente Orench, J. F. Clergeau, S. Martínez, M. Olmos, O. Fabelo, and J. Campo, The new powder diffractometer D1B of the Institut Laue Langevin, J. Phys.: Conf. Ser. 549, 012003 (2014).
- Parui et al. [2025] K. K. Parui, A. A. Kulbakov, E. Häußler, N. S. Pavlovskii, A. Mannathanath Chakkingal, M. Avdeev, R. Gumeniuk, S. Granovsky, A. Mistonov, S. A. Zvyagin, T. Doert, D. S. Inosov, and D. C. Peets, Disordered ground state in the 3D face-centred frustrated spin- system MnSn(OH) (2025), arXiv:2502.12433 [cond-mat.str-el] .
- Kulbakov et al. [2025b] A. A. Kulbakov, E. Häußler, A. Mannathanath Chakkingal, N. S. Pavlovskii, K. K. Parui, S. A. Granovsky, S. Gaß, L. T. Corredor Bohórquez, A. U. B. Wolter, S. A. Zvyagin, Y. V. Skourski, V. Y. Pomjakushin, I. Puente-Orench, D. C. Peets, T. Doert, and D. S. Inosov, Data underpinning: Spin liquid mimicry in the hydroxide double perovskite CuSn(OD)6 induced by correlated proton disorder (2025b), OPARA repository, Technische Universität Dresden, doi:10.25532/OPARA-790.
- [48] D. C. Peets, D. Inosov, A. Kulbakov, K. Parui, I. Puente Orench, and H. Yoshida, Magnetic structure of kagome In-kapellasite, InCu(OD)Cl, Institut Laue-Langevin (ILL), doi:10.5291/ILL-DATA.5-31-2952.